Circular No. 5908

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET WISEMAN-SKIFF (1993u) Suspected weak images of this comet on CCD frames obtained by B.

Schmidt with the Multiple-Mirror Telescope at Mt. Hopkins on Feb. 2 (measured by G. V. Williams and Schmidt) are confirmed with an independent recovery by J. V. Scotti in Spacewatch images obtained by T. Gehrels at Kitt Peak on Dec. 16 (low altitude, twilight, rather poor seeing), after which Scotti found the comet on a single frame obtained on Dec. 10. On Dec. 16, the comet showed a coma 13′′ in diameter (m_{2} = 22.6), with a faint 0′.34 tail in p.a. 286^{o}. The indicated correction to the prediction on MPC 16380 (ephemerides on MPC 20654 and 22512) is ΔT = −0.08 day.

1993 UT α_{2000} δ_{2000} m_{1} Observer
Feb. 2.08400 23^{h}22^{m}07.^{s}26 +10^{o}34′32′′.0  Schmidt
2.09154 23 22 08.00 +10 34 35.4  _{”}
2.09866 23 22 09.06 +10 34 42.2  _{”}
Dec. 10.52020 12 32 25.30 −21 05 46.8 19.6 Scotti
16.51914 12 39 23.51 −22 24 11.0  _{”}
16.52650 12 39 24.15 −22 24 16.3 20.8  _{”}
16.53339 12 39 24.42 −22 24 21.5  _{”}
16.54100 12 39 24.91 −22 24 25.9  _{”}
16.54850 12 39 25.59 −22 24 32.6  _{”}

PERIODIC COMET KUSHIDA-MURAMATSU (1993t) Numerous additional observations are available on MPEC 1993-Y02,

as are the following elliptical orbital elements, by S. Nakano, Sumoto, from 35 observations Dec. 8-15:

T = 1993 Nov. 4.459 TT  ω = 333.^{o}504
e = 0.35592  Ω = 96.737 }2000.0
q = 2.33989 AU  i = 1.673
a = 3.63294 AU n^{o} = 0.142337 P = 6.92 years
1993/94 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 9 5^{h}35.^{m}54 +22^{o}33.^{′}3 1.371 2.351 172.^{o}6 3.^{o}1 15.4
19 5 27.47 +22 37.0 1.377 2.359 175.3  2.0 15.4
29 5 20.03 +22 40.3 1.409 2.368 163.4  6.8 15.5
Jan. 8 5 14.39 +22 44.3 1.466 2.379 151.9 11.2 15.6
18 5 11.38 +22 50.1 1.546 2.392 141.1 15.0 15.7

1993 December 16 Daniel W. E. Green